Grade 10Physical Chemistry

How many millilitres of 0.1M H2SO4 must be added to 100 ml of 0.05M NaOH to give a solution that a concentration of 0.05M H2SO4?Pls tell it is urgent

tanishq chawla
9 Years agoGrade 10

1 Answer

varun
ApprovedApproved Tutor Answer9 Years ago
HELLO TANISHQ,
I Will give you the solution below see it
2 X 0.1V – 100 X 0.5 = 2 X 0.05(V X 100)
0.2V  – 5 = 0.1(V + 100)
0.2V – 5 = 0.1V + 10
=> V = 150ml
THANK YOU